# Multi-agent Adaptive Mechanism Design

### v1

@@ Abstract.
We study a sequential mechanism design problem in which a principal seeks to elicit truthful reports from multiple rational agents while starting with no prior knowledge of agents’ beliefs. We introduce Distributionally Robust Adaptive Mechanism (DRAM), a general framework combining insights from both mechanism design and online learning to jointly address truthfulness and cost-optimality. Throughout the sequential game, the mechanism would estimate agents’ beliefs, then iteratively updates a distributionally robust linear program with shrinking ambiguity sets to reduce payments while preserving truthfulness. Our mechanism guarantees truthful reporting with high probability while achieving $\tilde{O}(\sqrt{T})$ cumulative regret, and we establish a matching lower bound showing that no truthful adaptive mechanism can asymptotically do better. The framework generalizes to plug-in estimators ($\texttt{DRAM}+$), supporting structured priors and delayed feedback. To our knowledge, this is the first adaptive mechanism under the general settings that maintains truthfulness and achieves optimal regret when incentive constraints are unknown and must be learned.
@@ 1. Introduction
The theory of mechanism design studies rules and institutions in various disciplines, ranging from auctions and online advertisements to business contracts and trading rules. The formulation often involves a central principal (system) and one or many rational agents (players), where the principal designs a mechanism to achieve a given objective subject to agents’ incentives. A typical component in such formulations is the common knowledge assumption: certain information about agents is presumed known to the principal and can be exploited to design analytically tractable, often optimal mechanisms. For example, the design of revenue-optimal auctions requires knowledge of bidders’ value distributions over the item (myerson1981optimal). However, the availability of such knowledge is difficult to justify in practice. This observation, originally due to (wilson1985game), is now known as Wilson’s critique. It proposes that some information is too private to be common knowledge, and such assumptions should be weakened to approximate reality.
@@
In parallel, the theory of online learning studies algorithms that learn and make decisions in unfamiliar environments, aiming to approach the performance of oracles that have full knowledge from the start. The principal typically begins with no knowledge of the environment, and information is acquired through repeated data collection and carefully designed statistical methods. A common assumption is that the environment is unknown but stationary. For example, in the classical multi-armed bandit model (lattimore2020bandit), each arm’s reward is a stochastic distribution, and the best arm can be discovered via repeated sampling. An alternative is to assume the worst-case scenario from the environment, i.e., fully adversarial feedback. These algorithms have wide applications in recommendation, pricing, scheduling, and more (lattimore2020bandit). In application, however, they often interact with humans, who are neither stationary nor fully adversarial. In fact, a foundational assumption in economics is that humans are rational (von2007theory).
@@
Therefore, the strength and weaknesses from both fields seems to be complementary to each one. Mechanism design can incentivize nice behavior for proper learning guarantees from rational agents, while online learning can provide the necessary knowledge for efficient mechanisms. For this reason, the combination of mechanism design and online learning has received increasing attention, most notably in settings such as online contract design (ho2014adaptive; zhu2022sample) and online auctions (blum2004online; cesa2014regret). However, the design of general multi-agent adaptive mechanisms remains an under-explored problem.
@@
In this work, we consider a general sequential mechanism design problem in which the principal starts with no knowledge of the players. As a motivating example, consider the image labeling task where the principal assigns unlabeled images to multiple agents. In each round, each agent makes an independent observation of the image as her type, then reports her observation (type), and finally receives a payment from the principal. The principal’s objective is three-fold: data quality, truthfulness, and cost-optimality. The principal wants to design a reward mechanism that can obtain the highest-quality data from the crowdsourced task, while incentivizing truthful report from players, and do so in a cost-minimal way.
@@
As the principal cannot directly and cheaply observe ground truth at scale, the design of such a reward mechanism faces several challenges. First, the true label is unavailable or expensive to obtain for most tasks (otherwise the principal would prefer that route), making it hard to control report quality or infer agent skills. Second, each agent’s observation (type), or whatever actions she does before reporting, is unknown to the principal. Agents are rational and pursue utility, so they may lie or become lazy (skip observation and report random labels) during this phase. Finally, classical mechanism designs that rely on common knowledge are inapplicable: even mechanisms that only aim to elicit truthful reports often assume accurate knowledge of posteriors or correlation structure (miller2005eliciting). In addition, we cannot exploit the special structures in specific mechanism design problems (such as in auctions, a second-price auction incentivizes truthfulness even without common knowledge). For a general mechanism design problem, the principal must learn this knowledge while simultaneously incentivizing effort and honesty even when such knowledge is inaccurate.
@@ 1.1. Our Contribution
Our model can be viewed as a generalization of the optimal mechanism design problem in the mechanism design field and prediction with expert advice problem in the online learning field. From a mechanism design perspective, we relaxed the common knowledge assumption. The principal would rely on online learning techniques to gain knowledge. From an online learning perspective, we relaxed the setting from always-honest agents to the more realistic rational agents. To incentivize honest work, the principal would utilize proper mechanism design.
The necessity of truthfulness. We show that in our sequential setting, mechanisms that aim to maximize downstream decision quality while reducing payments over time must, up to label permutations, induce truthful reports under mild regularity conditions. This result clarifies the target of design: once reports are (approximately) truthful, the principal can consistently learn the distributions required by the payment rule and drive down payments. The statement is structural rather than the main focus, and it guides the robust and adaptive designs that follow.
Distributionally robust mechanisms. We formulate the single-round mechanism design problem as a linear program that minimizes expected payment subject to (i) individual rationality, (ii) truthfulness, and (iii) a no-free-lunch constraint for uninformative strategies. Because the principal initially misspecifies the required knowledge, we introduce distributional robustness: incentive constraints are enforced against all posteriors within a total-variation ambiguity set around current estimates. The resulting mechanism preserves truthfulness under bounded misspecification, at a quantifiable payment premium that scales with the ambiguity radius. As data accumulates, the radius shrinks and payments decrease accordingly.
Optimal adaptive mechanism design. Finally, in the sequential mechanism design problem, we show that the fundamental bottleneck lies in learning players’ beliefs. This learning problem is unavoidable: without sufficiently accurate knowledge of players’ beliefs, no mechanism can simultaneously guarantee truthfulness and low payment. Motivated by this observation, we design an algorithm named Distributionally Robust Adaptive Mechanism (DRAM) that carefully balances robustness and learning. Our algorithm achieves an $\tilde{O}(N\sqrt{T})$ regret guarantee (up to logarithmic factors) while preserving truthfulness with high probability at every round, and we complement this result with a matching lower bound showing that no truthful adaptive mechanism can do better in the worst case. The theoretical results are validated by numerical simulations. This framework extends to any plug-in estimators (e.g., structured or regularized estimators for discrete distributions) and is compatible with delayed or batched feedback. To our knowledge, this is the first general adaptive mechanism that maintains truthfulness and achieves optimal regret when incentive constraints depend on unknown and learned information.
@@ 1.2. Related Work
Our work draws insights from both the mechanism design and online learning literature, including adaptive mechanism design and online learning.
@@ 1.2.1. Online and adaptive Mechanism Design
The combination of mechanism design and online learning is a fast-growing direction in algorithmic game theory (roughgarden2010algorithmic). Prior work spans a variety of participation structures and problem settings. Some works study synchronous settings, where the same players have multiple encounters with each other, as in repeated games (hart2000simple; satchidanandan2023incentive; papadimitriou2022complexity). Others consider asynchronous settings, where new agents may arrive and depart over time, a structure particularly relevant in online auctions and advertising platforms (milgrom2019auction; choi2020online; hajiaghayi2004adaptive; cesa2007improved; wang2017display). Within both settings, two domains have received the most attention: online contract design (ho2014adaptive; zhu2022sample) and online auctions (blum2004online; cesa2014regret). These works typically focus on learning an optimal mechanism, such as an optimal contract or an optimal reserve price, using tools from bandit learning.
A key observation is that preserving truthfulness is not a substantive difficulty in these existing models. In contract design, agents’ choices naturally reflect their private information, and no truthful reporting constraint is involved. In online auctions, structural properties ensure incentive compatibility: for instance, in second-price auctions, truthful bidding remains a dominant strategy even when the reserve price is inaccurate. Even without any knowledge, the second-price auction guarantee’s players’ truthfulness. As a result, works such as (cesa2014regret) can safely explore suboptimal mechanisms during learning without risking incentive distortion or data contamination. The focus is solely on learning an optimal mechanism, which makes these problems are more or less reducible to a bandit problem (zhu2022sample; cesa2014regret).
In contrast, in a general mechanism design problem, preserving truthfulness becomes a significant difficulty. When the principal begins with limited or mis-specified knowledge, an improperly constructed mechanism can immediately encourage agents to lie or exert low effort, thereby corrupting the collected data undermining subsequent learning processes. Thus, unlike prior literature, maintaining truthfulness throughout the entire learning trajectory is not merely desirable but essential, and this requirement is one of the central challenge we must address.
@@ 1.2.2. Prediction with Expert Advice
We note that our setup is a stochastic, label-efficient variant of the classical prediction with expert advice problem. The prediction with expert advice problem is fundamental in online learning (cesa2006prediction). In the standard framework, the players (often called “experts”) can sequentially provides arbitrary and even adversarial signals, and the principal’s objective is to implement an aggregation algorithm that achieves sublinear decisional regret. A simplification is to assume players behave stochastically (report signals according to a probability law), under which the aggregation regret can be significantly improved (cesa2004generalization). The stochastic variant also has connections with other online learning problems such as online optimization (agarwal2017second; cesa2007improved; gaillard2014second), with extensions in contextual or non-stationary settings (besbes2016optimization).
In the practical setting, acquiring a true label might be expensive. It may be only feasible to query the true label for a small portion of rounds. This is the label-efficient setting of prediction with expert advice (helmbold1997some; cesa2005minimizing). Roughly speaking, the aggregation regret decreases as the inverse square root of the number of queries. There also exist adaptive algorithms that achieve the same regret with far fewer queries in benign cases (mitra2020adaptivity; castro2023adaptive).
@@ 1.2.3. Information elicitation and peer prediction
The field of information elicitation studies the mechanism design task to incentivize honest feedback from untrusted but rational participants, generally via designing scoring rules (li2022optimization) as rewards or penalties for participants. Particularly, peer prediction (miller2005eliciting) studies the scenarios in which ground truth is unavailable for direct verification of collected reports, with applications in dataset acquisition and evaluation (chen2020truthful; zheng2024proper), crowdsourcing (dasgupta2013crowdsourced), and recent blockchain-based decentralized ecosystems (wang2023uncertainty; zhao2024takes). The general paradigm of peer prediction mechanisms is to ask multiple participants the same question and reward them according to the comparison among their reports. While peer prediction mechanisms provide elegant results on truthful Nash equilibria without requirement of ground-truth information, most of existing mechanism rely on strong unrealistic assumptions on know prior and observation matrices, forming a gap to practical usage in real-world systems.
For practical usage, researcher develop a series of works with relaxed assumptions or stronger incentive guarantees. Particularly, (kong2024dominantly) develops a prior-free multi-task peer prediction mechanism with dominant-strategy incentive compatibility, with a lack of permutation-proof property that is impossible for any prior-free peer prediction mechanisms (kong2019information). Besides, (shnayder2016informed) provides an informed truthful mechanism ensuring that the truthful equilibrium achieves highest utilities among all Nash equilibria, and (zhang2025stochastically) develops a mechanism with a stochastic dominance property ensuring incentive compatibility even under non-linear utilities.
@@ 2. Problem Formulation
We consider the sequential mechanism design problem where a principal seeks to elicit truthful reports from rational agents. The principal sequentially assigns $T$ prediction tasks to a group of $N$ rational players. Each task has a true label $Y_{t}\in\mathcal{Y}$, i.i.d. sampled from an unknown and stationary distribution $p_{Y}(\cdot)$. Unless stated otherwise, this true label is not revealed to anyone, either the principal or the players. We let $\mathcal{Y}$ be finite to avoid mathematical complications, and assume each true label $y$ appears with uniformly bounded probability $\underline{p}\leq p_{Y}(y)\leq\overline{p}$. In each round, each player $i=1,\dotsb,n$ independently studies the task, acquiring her own observation $X_{it}\in\mathcal{Y}$ with a constant cost $c$. $X_{it}$ is generated according to the player’s skill $p_{i}(x\mid y)$, a stationary conditional probability law. We assume $p_{i}$ is non-degenerate, i.e., there exists two labels $y,y^{\prime}$ where $p_{i}(\cdot\mid y)\neq p_{i}(\cdot\mid y^{\prime})$. In other words, observation should bring new information by stochastically distinguishing at least two of the labels. Each player might know her own skill distribution $p_{i}$, but has no information about anyone else’s, and the principal initially knows none of them. Aside from observation, players also has an outside option of lazily reporting a random label without observing the label (does not incur cost $c$ as well).
After studying the task, players independently produce their public reports $Z_{it}\in\mathcal{Y}$ to the principal. We assume players are risk-neutral and myopic. Being risk-neutral means players aim to maximize their expected reward, conditional on the public and private information they have. Being myopic means players only care about immediate reward in the current round but not future rewards. Under these settings, we have rational players who do not necessarily report their observations. Instead, they would lie (report $Z_{it}\neq X_{it}$) or be lazy (report $Z_{it}$ without observation $X_{it}$) when they expect an advantage in doing so. We denote the observation and report profile of all players in a round by $\boldsymbol{X}_{t}$ and $\boldsymbol{Z}_{t}$.
Collecting the reports, the principal rewards each player $i$ with an reward mechanism $R_{it}(\boldsymbol{Z}_{1},\dotsb,\boldsymbol{Z}_{t})$. The reports can then be used for downstream decision-making tasks, such as aggregation. Note that the reward mechanism is non-anticipating, meaning the mechanism can only decide on past and current but not future report profiles.
@@
  • Truthfulness (aka incentive-compatibility): given all other players act honestly, a player would maximize her own expected utility when she works, obtains observations, and then reports honestly ($Z_{it}=X_{it}$).
@@
  • Cost-optimality: maintaining truthfulness and data-quality, the principal minimizes its total expected payment to players.
We now compare between our setup and typical modeling assumptions in the online learning and mechanism design literature. Online learning mainly targets statistical efficiency—minimizing cumulative prediction error—while treating all reports as truthful. Mechanism design, by contrast, centers on strategic incentives, but usually presumes players’ type distributions are known or even common knowledge. These assumptions ease analysis, yet rarely hold in practice. Our model pursues both goals at once and relaxes the assumptions from both fields.
@@
The proposed model can be further generalized to match the classical model in the mechanism design literature. Here, players’ observations are their own types of the round. Assume in each round players’ types $\boldsymbol{X}$ are sampled from a stationary joint distribution. Players then report their types (not necessarily truthful) $\boldsymbol{Z}$ to the principal and receive rewards. All of our analysis and algorithms applies to this generalized setting. In fact, our analysis does not require how the types are generated and make no use of $p_{Y}$ and $p_{i}$ and focuses exclusively on the joint law $p_{\boldsymbol{X}}$.
@@
We note that each agent’s utility is linear in only her own reward ($u_{i}=r_{i}$). In the most general setting of mechanism design, an agent’s utility is a function of all agents’ types and the resource allocation from principal: $u_{i}:\mathcal{Y}^{N}\times\mathcal{R}\to\mathbb{R}$, where $\mathcal{R}$ is the space of the principal’s resource allocation decisions. For example, in contracts, utility depends on the agent’s own type $x_{i}$ and the principal’s payment $r_{i}$ ($u_{i}=f(x_{i},r_{i})$). In auctions, utility depends on whether or not the agent gets the item (with probability $p$), her valuation of the item (type $x_{i}$), and her requested payment $r_{i}$ ($u_{i}=p\cdot x_{i}-r_{i}$). Our analysis could be potentially generalized to the case when such utility function is exactly known by the principal. Nonetheless, we believe substantial hurdle exists when such utility functions are not known and need to be learned.
We conclude this section by revealing the importance of truthfulness. After all, the principal’s top objectives in outsourcing tasks are to improve data quality and lower costs. Truthfulness, as a mechanism design objective, might not be of interest if the mechanism that reaches the highest quality or the lowest cost promotes dishonest behaviors. From the revelation principle (myerson1979incentive), we know that truthfulness is “free”, in the sense that we don’t lose anything by focusing only on mechanisms with their incentive-compatible Nash equilibria. For the same reason, it suffices to consider the setting where the true label $Y$, observation $X$, and report $Z$ all belong to the same space $\mathcal{Y}$. The following proposition actually proves a stronger result, showing that truthfulness is not only free, but in fact almost necessary for maximal quality.
@@
Consider a general non-anticipating decision-making task $A_{t}(\boldsymbol{Z}_{1},\dotsb,\boldsymbol{Z}_{t})$ with objective $\operatorname{Obj}(A_{t},Y_{t})$. Assume players’ skills $p_{i}$ are non-degenerate. For any round $t$, the highest possible performance is attainable if and only if every player first observes the task, and then either

### v2

@@ Abstract.
We study the sequential mechanism design problem in which a principal seeks to elicit truthful reports from multiple rational agents while starting with no prior knowledge of agents’ beliefs. We introduce Distributionally Robust Adaptive Mechanism (DRAM), a general framework combining insights from both mechanism design and online learning to jointly address truthfulness and cost-optimality. Throughout the sequential game, the mechanism would estimate agents’ beliefs, then iteratively updates a distributionally robust linear program with shrinking ambiguity sets to reduce payments while preserving truthfulness. Our mechanism guarantees truthful reporting with high probability while achieving $\tilde{O}(N\sqrt{T})$ cumulative regret, and we establish a matching lower bound showing that no truthful adaptive mechanism can asymptotically do better. The framework generalizes to plug-in estimators ($\texttt{DRAM}+$), supporting structured priors and delayed feedback. To our knowledge, this is the first adaptive mechanism under the general settings that maintains truthfulness and achieves optimal regret when incentive constraints are unknown and must be learned.
@@ 1. Introduction
The theory of mechanism design studies rules and institutions in various disciplines, ranging from auctions and online advertisements to business contracts and trading rules. The formulation often involves a central principal (system) and one or many rational agents (players), where the principal designs a mechanism to achieve a given objective subject to agents’ incentives. A typical component is the common knowledge assumption: certain information about agents is presumed known to the principal and can be exploited to design analytically tractable, often optimal mechanisms. For example, knowledge over bidders’ value distributions over the auctioned can be used to design revenue-optimal auctions (Myerson, 1981). However, the availability of such knowledge is difficult to justify in practice. This observation, originally due to (Wilson, 1985), is now known as Wilson’s critique. It proposes that some information is too private to be common knowledge, and such assumptions should be weakened to approximate reality.
@@
In parallel, the theory of online learning studies algorithms that learn and make decisions in unfamiliar environments, aiming to approach the performance of oracles that have full knowledge from the start. The principal typically begins with no knowledge of the environment, and information is acquired through repeated data collection and carefully designed statistical methods. A common assumption is that the environment is unknown but stationary. For example, in the classical multi-armed bandit model (Lattimore and Szepesvári, 2020), each arm’s reward is a stochastic distribution, and the best arm can be discovered via repeated sampling. An alternative is to assume the worst-case scenario from the environment, i.e., fully adversarial feedback. These algorithms have wide applications in recommendation, pricing, scheduling, and more (Lattimore and Szepesvári, 2020). In application, however, they often interact with humans, who are neither stationary nor fully adversarial. In fact, a foundational assumption in economics is that humans are rational (Von Neumann and Morgenstern, 2007).
@@
Therefore, the strength and weaknesses from both fields seems to complement to each one. Mechanism design incentivizes nice behavior from rational agents for proper learning guarantees, and online learning can provide the necessary knowledge for efficient mechanisms. For this reason, the combination of mechanism design and online learning has received increasing attention, most notably in settings such as online contract design (Ho et al., 2014; Zhu et al., 2022) and online auctions (Blum et al., 2004; Cesa-Bianchi et al., 2014). However, the design of general multi-agent adaptive mechanisms remains an under-explored problem.
@@
In this work, we study the sequential mechanism design problem in which a principal in each round designs reward mechanisms for multiple rational agents, while starting with no prior knowledge of agents’ beliefs. The principal’s objective is three-fold: data quality, truthfulness, and cost-optimality. The principal wants to design a reward mechanism that can obtain the highest-quality data from the task, while incentivizing truthful report from agents, and do so in a cost-minimal way. As a motivating example, consider the image labeling task where the principal assigns raw images to multiple agents for labeling (Figure 1). In each round, each agent makes a private observation of the image as her type, then reports her type to the principal, and finally receives a payment.
@@
The multi-agent mechanism design problem faces several major challenges. First, each agent’s observation is private and unknown to the principal. Agents are rational and pursue utility, so they may lie or become lazy. However, ground truth may be unavailable or expensive to obtain, making it hard to directly control report quality or infer agent skills. Second, classical mechanism designs that rely on common knowledge are inapplicable: even mechanisms that only focus on maintaining truthfulness often assume accurate knowledge of posteriors or correlation structure (Miller et al., 2005). Finally, we cannot exploit the structures in specific mechanism design problems. For example, in auction design, a second-price auction incentivizes truthfulness without any common knowledge. However, for a general mechanism design problem, maintaining truthfulness itself under limited knowledge is already a substantial difficulty.
@@ 1.1. Our Contribution
Our work draws insights from both the mechanism design and online learning literature. From a mechanism design perspective, we relaxed the common knowledge assumption, and generalizes the optimal mechanism design problem. From an online learning perspective, we relaxed the setting from always-honest agents to the more realistic rational agents, and generalizes the prediction with expert advice problem.
The necessity of truthfulness. We show that truthfulness is in a sense “necessary” for sequential decision making. Any decision making process based on agents’ rational reports can achieve highest performance if and only if agents are truthful (up to permutations). This is a result based on Blackwell’s informativeness theorem (Blackwell, 1953), and stronger than the revelation principle (Myerson, 1979). Since mechanism design is also a decision making task itself, it implies truthfulness is necessary for learning optimal mechanisms.
Distributionally robust mechanisms. We introduce and study a family of distributionally robust mechanisms, which preserves truthfulness and incurs low cost even when the principal’s knowledge is ambiguous. We study the relations between design parameters and achievable robustness, establish methods to tractably acquire these mechanisms, and finally characterize the cost of robustness.
Optimal adaptive mechanism design. We design a general framework named Distributionally Robust Adaptive Mechanism (DRAM), which estimate agents’ beliefs and iteratively updates a distributionally robust mechanism based on estimation accuracy. Our algorithm achieves an $\tilde{O}(N\sqrt{T})$ regret guarantee (up to logarithmic factors) while preserving truthfulness with high probability. We complement this result with a matching lower bound showing that no truthful adaptive mechanism can do better in the worst case. The theoretical results are validated by numerical simulations. This framework extends to any plug-in estimators (e.g., structured or regularized estimators for discrete distributions) and is compatible with delayed or batched feedback. To our knowledge, this is the first general adaptive mechanism that maintains truthfulness and achieves optimal regret when incentive constraints depend on unknown and learned information.
@@ 1.2. Related Work
Our work draws insights from both the mechanism design and online learning literature.
@@ 1.2.1. Online and adaptive Mechanism Design
The combination of mechanism design and online learning is a fast-growing direction in algorithmic game theory (Roughgarden, 2010). Prior work spans a variety of participation structures and problem settings. Some works study synchronous settings, where the same agents have multiple encounters with each other, as in repeated games (Hart and Mas-Colell, 2000; Satchidanandan and Dahleh, 2023; Papadimitriou et al., 2022). Others consider asynchronous settings, where new agents may arrive and depart over time, a structure particularly relevant in online auctions and advertising platforms (Milgrom, 2019; Choi et al., 2020; Hajiaghayi et al., 2004; Cesa-Bianchi et al., 2007; Wang et al., 2017). Under both settings, two domains have received the most attention: online contract design (Ho et al., 2014; Zhu et al., 2022) and online auctions (Blum et al., 2004; Cesa-Bianchi et al., 2014). These works typically focus on learning an optimal mechanism, such as an optimal contract or an optimal reserve price, using tools from bandit learning.
A key observation is that preserving truthfulness is not a substantive difficulty in these existing models. In contract design, agents’ choices naturally reflect their private information, and no truthful reporting constraint is involved. In online auctions, structural properties ensure incentive compatibility: for instance, in second-price auctions, truthful bidding remains a dominant strategy even when the reserve price is inaccurate. Even without any knowledge, the second-price auction guarantee’s agents’ truthfulness. As a result, works such as (Cesa-Bianchi et al., 2014) can safely explore suboptimal mechanisms during learning without risking incentive distortion or data contamination. The focus is solely on learning an optimal mechanism, which makes these problems are more or less reducible to a bandit problem (Zhu et al., 2022; Cesa-Bianchi et al., 2014).
In contrast, in a general mechanism design problem, preserving truthfulness becomes a significant difficulty. When the principal begins with ambiguous knowledge, an improperly constructed mechanism can immediately encourage agents to lie or exert low effort, thereby corrupting the collected data, undermining subsequent learning processes. Thus, unlike prior literature, maintaining truthfulness throughout the entire learning trajectory is not merely desirable but essential, and this requirement is one of the central challenge we must address.
@@ 1.2.2. Prediction with Expert Advice
We note that our setup is a stochastic, label-efficient variant of the classical prediction with expert advice problem. The prediction with expert advice problem is fundamental in online learning (Cesa-Bianchi and Lugosi, 2006). In the standard framework, the agents (often called “experts”) can sequentially provides arbitrary and even adversarial signals, and the principal’s objective is to implement an aggregation algorithm that achieves sublinear regret. A simplification is to assume agents behave stochastically (report signals according to a probability law), under which the aggregation regret can be significantly improved (Cesa-Bianchi et al., 2004). The stochastic variant also has connections with other online learning problems such as online optimization (Agarwal et al., 2017; Cesa-Bianchi et al., 2007; Gaillard et al., 2014), with extensions in contextual or non-stationary settings (Besbes et al., 2016).
In the practical setting, acquiring a true label might be expensive. It may be only feasible to query the true label for a small portion of rounds. This is the label-efficient setting of prediction with expert advice (Helmbold and Panizza, 1997; Cesa-Bianchi et al., 2005). Roughly speaking, the aggregation regret decreases as the inverse square root of the number of queries. There also exist adaptive algorithms that achieve the same regret with far fewer queries in benign cases (Mitra and Gopalan, 2020; Castro et al., 2023).
@@ 1.2.3. Information elicitation and peer prediction
The field of information elicitation studies the mechanism design task to incentivize honest feedback from untrusted but rational participants, generally via designing scoring rules (Li et al., 2022) as rewards or penalties for participants. Particularly, peer prediction (Miller et al., 2005) studies the scenarios in which ground truth is unavailable for direct verification of collected reports, with applications in dataset acquisition and evaluation (Chen et al., 2020; Zheng et al., 2024), crowdsourcing (Dasgupta and Ghosh, 2013), and recent blockchain-based decentralized ecosystems (Wang et al., 2023; Zhao et al., 2024). The general paradigm of peer prediction mechanisms is to ask multiple participants the same question and reward them according to the comparison among their reports. While peer prediction mechanisms provide elegant results on truthful Nash equilibria without requirement of ground-truth information, most of existing mechanism rely on strong unrealistic assumptions on know prior and observation matrices, forming a gap to practical usage in real-world systems.
For practical usage, researcher develop a series of works with relaxed assumptions or stronger incentive guarantees. Particularly, (Kong, 2024) develops a prior-free multi-task peer prediction mechanism with dominant-strategy incentive compatibility, with a lack of permutation-proof property that is impossible for any prior-free peer prediction mechanisms (Kong and Schoenebeck, 2019). Besides, (Shnayder et al., 2016) provides an informed truthful mechanism ensuring that the truthful equilibrium achieves highest utilities among all Nash equilibria, and (Zhang et al., 2025) develops a mechanism with a stochastic dominance property ensuring incentive compatibility even under non-linear utilities.
@@ 2. Problem Formulation
We consider the sequential mechanism design problem where a principal seeks to elicit truthful reports from rational agents. The principal sequentially assigns $T$ prediction tasks to a group of $N$ rational agents. Each task has a true label $Y_{t}\in\mathcal{Y}$, i.i.d. sampled from an unknown and stationary distribution $p_{Y}(\cdot)$. Unless stated otherwise, this true label is not revealed to anyone, either the principal or the agents. We let $\mathcal{Y}$ be finite to avoid mathematical complications, and assume each true label $y$ appears with uniformly bounded probability $\underline{p}\leq p_{Y}(y)\leq\overline{p}$. In each round, each agent $i=1,\dotsb,n$ independently studies the task, acquiring her own observation $X_{it}\in\mathcal{Y}$ with a constant cost $c$. $X_{it}$ is generated according to the agent’s skill $p_{i}(x\mid y)$, a stationary conditional probability law. We assume $p_{i}$ is non-degenerate, i.e., there exists two labels $y,y^{\prime}$ where $p_{i}(\cdot\mid y)\neq p_{i}(\cdot\mid y^{\prime})$. In other words, observation should bring new information by stochastically distinguishing at least two of the labels. Each agent might know her own skill distribution $p_{i}$, but has no information about anyone else’s, and the principal initially knows none of them. Aside from observation, agents also has an outside option of lazily reporting a random label without observing the label (does not incur cost $c$ as well).
After studying the task, agents independently produce their public reports $Z_{it}\in\mathcal{Y}$ to the principal. We assume agents are risk-neutral and myopic. Being risk-neutral means agents aim to maximize their expected reward, conditional on the public and private information they have. Being myopic means agents only care about immediate reward in the current round but not future rewards. Under these settings, we have rational agents who do not necessarily report their observations. Instead, they would lie (report $Z_{it}\neq X_{it}$) or be lazy (report $Z_{it}$ without observation $X_{it}$) when they expect an advantage in doing so. We denote the observation and report profile of all agents in a round by $\boldsymbol{X}_{t}$ and $\boldsymbol{Z}_{t}$. Note that $p_{Y}$ and $p_{i}$ together defines a joint law $p_{\boldsymbol{X}}$ over $\boldsymbol{X}\in\mathcal{Y}^{N}$, and we later show that learning this $p_{\boldsymbol{X}}$ is crucial for optimal mechanisms.
Collecting the reports, the principal rewards each agent $i$ with an reward mechanism $R_{it}(\boldsymbol{Z}_{1},\dotsb,\boldsymbol{Z}_{t})$. The reports can then be used for downstream decision-making tasks, such as aggregation. Note that the reward mechanism is non-anticipating, meaning the mechanism can only decide on past and current but not future report profiles.
@@
  • Truthfulness (aka incentive-compatibility): given all other agents act honestly, a agent would maximize her own expected utility when she works, obtains observations, and then reports honestly ($Z_{it}=X_{it}$).
@@
  • Cost-optimality: maintaining truthfulness and data-quality, the principal minimizes its total expected payment to agents.
We now compare between our setup and typical modeling assumptions in the online learning and mechanism design literature. Online learning mainly targets at minimizing cumulative decisional error, while treating all reports as truthful. Mechanism design, by contrast, centers on strategic incentives, but usually presumes agents’ type distributions are known or even common knowledge. These assumptions ease analysis, yet rarely hold in practice. Our model pursues both goals at once and relaxes the assumptions from both fields.
@@
The proposed model can be further generalized to match the classical model in the mechanism design literature. Here, agents’ observations are their own types of the round. Assume in each round agents’ types $\boldsymbol{X}$ are sampled from a stationary joint distribution. agents then report their types (not necessarily truthful) $\boldsymbol{Z}$ to the principal and receive rewards. All of our analysis and algorithms applies to this generalized setting. In fact, our analysis does not require how the types are generated and make no use of $p_{Y}$ and $p_{i}$ and focuses exclusively on the joint law $p_{\boldsymbol{X}}$.
@@
We note that each agent’s utility is linear in only her own reward ($u_{i}=r_{i}$). In the most general setting of mechanism design, an agent’s utility is a function of all agents’ types and the resource allocation from principal: $u_{i}:\mathcal{Y}^{N}\times\mathcal{R}\to\mathbb{R}$, where $\mathcal{R}$ is the space of the principal’s resource allocation decisions. For example, in contracts, utility depends on the agent’s own type $x_{i}$ and the principal’s payment $r_{i}$ ($u_{i}=f(x_{i},r_{i})$). In auctions, utility depends on whether or not the agent gets the item (with probability $p$), her valuation of the item (type $x_{i}$), and her requested payment $r_{i}$ ($u_{i}=p\cdot x_{i}-r_{i}$). Our analysis may potentially be generalized to the case when such utility function is exactly known by the principal.
We conclude this section by revealing the importance of truthfulness. After all, the principal’s top objectives in outsourcing tasks are to improve data quality and lower costs. Truthfulness, as a mechanism design objective, might not be of interest if the mechanism that reaches the highest quality or the lowest cost promotes dishonest behaviors. From the revelation principle (Myerson, 1979), we know that truthfulness is “free”, in the sense that we don’t lose anything by focusing only on mechanisms with their incentive-compatible Nash equilibria. For the same reason, it suffices to consider the setting where the true label $Y$, observation $X$, and report $Z$ all belong to the same space $\mathcal{Y}$. The following proposition actually proves a stronger result, showing that truthfulness is not only free, but in fact almost necessary for maximal quality.
@@
Consider a general non-anticipating decision-making task $A_{t}(\boldsymbol{Z}_{1},\dotsb,\boldsymbol{Z}_{t})$ with objective $\operatorname{Obj}(A_{t},Y_{t})$. Assume agents’ skills $p_{i}$ are non-degenerate. For any round $t$, the highest possible performance is attainable if and only if every agent first observes the task, and then either